# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502 Japan

### v2

@@ Appendix E Finite size effect on the correlation function
$xx$ 1.12(3) 0.107(5) 1.32 $[4,10]$
0 $yy$ 1.07(3) 0.098(4) 0.66 $[4,10]$
$zz$ 1.13(4) 0.107(6) 0.25 $[4,10]$
$xx$ 0.35(6) 0.0149(19) 0.99 $[7,16]$
8 $yy$ 0.22(3) 0.0100(6) 0.49 $[4,18]$
$zz$ 0.20(4) 0.0095(7) 0.38 $[4,18]$
@@
In the case of $\lambda=8$, the best-fit values of $p$ at $L=40$ in Table 3 are larger than those at $L=32$ in Table 6. Thus, if we take the thermodynamic limit $L\to\infty$, there is a chance to get extrapolated values satisfying $p>1$, which corresponds to a positive value of the critical exponent $\eta$. Although we mentioned the inconsistency between the values of $\eta$ estimated from the magnetic susceptibility and correlation function in Section 6, it might be resolved by taking the thermodynamic limit in future research.
Appendix F Correlation function at off-peak temperatures

### v3

@@ Appendix F Finite size effect on the correlation function
xxxx 1.12(3) 0.107(5) 1.32 $[4,10]$
0 yyyy 1.07(3) 0.098(4) 0.66 $[4,10]$
zzzz 1.13(4) 0.107(6) 0.25 $[4,10]$
xxxx 0.35(6) 0.0149(19) 0.99 $[7,16]$
8 yyyy 0.22(3) 0.0100(6) 0.49 $[4,18]$
zzzz 0.20(4) 0.0095(7) 0.38 $[4,18]$
@@
In the case of $\lambda=8$, the best-fit values of $p$ at $L=40$ in Table 3 are larger than those at $L=32$ in Table 6. Thus, if we take the thermodynamic limit $L\to\infty$, there is a chance to get extrapolated values satisfying $p>1$, which corresponds to a positive value of the critical exponent $\eta$. Although we mentioned the inconsistency between the values of $\eta$ estimated from the magnetic susceptibility and correlation function in Section VI, it might be resolved by taking the thermodynamic limit in future research.
Appendix G Correlation function at off-peak temperatures